OpenStudy (anonymous):

Solve the equation. Check the solution. -2/x+4 = 4/x+3 PLEASE HELP ME PLEASE ILL BECOME A FN BEST ANSWER WHATEVER BUT PLEASE HURRY

Directrix (directrix):

@greenandblue Is the problem -2/(x+4) = 4/(x+3) OR (-2/x) + 4 = (4/x) + 3

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

\[\frac{-2}{x+4}=\frac{4}{x+3}\] \[-2(x+3)=4(x+4)\] \[-2x+6=4x+16\] \[6=6x+16\] etc

OpenStudy (anonymous):

and the cross mulitple i tried that but none of the choices are there for it....

OpenStudy (anonymous):

then just finish it from there ...like subtract 16 from 6 and stuff right?

OpenStudy (anonymous):

thanks @Directrix :)

OpenStudy (anonymous):

and everyone:)